Each of eighteen students was given a different integer from 1 through 18, inclusive. When asked to select a dance partner, the sum of the numbers for each of the nine couples was a perfect square. Who was dancing with whom?
Answer by checkley77(12844)

NUMBERS DANCING TOGETHER ADD UP TO 16 OR 25 BOTH SQUARES.
18+7=25
17+8=25
16+9=25
15+1=16
14+2=16
13+3=16
12+4=16
11+5=16
10+6=16